# Uncertainty modeling for fine-tuned implicit functions

### v1

@@ 1 Introduction
Recent advancements in neural implicit functions have facilitated their use for 3D object representations and applications in novel views synthesis in computer vision. Neural Radiance Fields (NeRFs) [1] have gained recognition for their ability to accurately synthesize novel views of complex scenes and became an important tool in applications of photorealistic rendering [2], such as virtual reality and augmented reality. Signed Distance Functions (SDFs)[3, 4] is shown to be particularly useful in scenarios where precise boundary details are crucial, like industrial design and robotics. Occupancy Networks[5], which model shapes as a probabilistic grid of space occupancy, excel in handling topological variations, making them ideal for medical imaging[6] and animation. Advancing this concept, Convolutional Occupancy Networks [7] integrate convolutional networks to enhance spatial learning, proving effective in detailed architectural modeling and complex reconstructions.
@@
In practice, achieving optimal performance with these methods often requires densely sampled input data and a high degree of similarity between the training data and the target object. However, such ideal conditions are rarely met in practical applications like augmented reality, virtual reality, autonomous driving, and medical contexts, where inputs are typically sparse and less precise[8, 9]. For example, in medical applications, generating precise 3D representations from sparse inputs is particularly crucial for morphological analysis[10]. Patient-specific anatomical modeling significantly enhances the assessment of a patient’s condition and aids in devising customized treatment plans[11].
@@
In addition to sparsity, real-world data usually suffers from noise and corruption that leads to distribution shifts with respect to the training data, and thus to significant performance degradation. Occlusion, noise, truncation, and lack of depth measurements [12] greatly affect reconstruction from monocular observations. Noise in estimations of camera poses inevitably degrades reconstruction from sparse views[8, 13, 14]. Under input sparsity and distribution shifts, in safety-critical applications, such as medical[15] or autonomous driving[16], it is crucial that inferred information is transparently disclosed to the end user, as it may significantly influence the decision-making process. One approach to this end is to quantify uncertainty in the reconstruction. However, very few works focused on modeling uncertainty in these implicit representations[12].
This paper addresses uncertainty quantification in 3D reconstruction from sparse and corrupted data using neural implicit functions. We propose Dropsembles, [^1] https://github.com/klanita/Dropsembles a method that leverages a noise-free, densely sampled dataset to model a shape prior and reconstructs based on sparse and noisy input through fine-tuning the model. Our approach aims to utilize the high-quality prior dataset to fill gaps and correct corruptions in the target data despite distribution shifts between prior and target datasets induced by noise and corruption. Crucially, the model is designed to recognize and highlight areas of high uncertainty. To the best of our knowledge, uncertainty quantification in 3D reconstruction for neural implicit functions has not been comprehensively addressed in the literature.
@@
Monte Carlo (MC) dropout [17] and deep ensembles [18] are two commonly used baselines for estimating uncertainty in computer vision applications [19], which could be readily applied to our task. MC dropout is simple to integrate and computationally efficient during training. However, it often underestimates uncertainty and requires multiple forward passes during inference, increasing computational cost[20, 21]. Deep ensembles involve training multiple neural networks independently and averaging their predictions, capturing a wider range of potential outputs. This method provides improved predictive performance and better-calibrated uncertainty estimates but is computationally costly and memory-intensive. This becomes especially demanding in our fine-tuning context, where each model in an ensemble needs to be trained on both datasets. Here we introduce Dropsembles, a method that creates ensembles based on the dropout technique. Dropsembles aim to moderate the computational demands associated with ensembles while attempting to maintain prediction accuracy and uncertainty calibration of deep ensembles. Combined with Elastic Weight Consolidation (EWC) [22], Dropsembles is able to mitigate distribution shifts between source and target datasets.
@@
  • We introduce Dropsembles (overview in Figure 1) to achieve the performance of vanilla ensembles with significantly reduced computation cost (subsection 4.1).
  • We introduce EWC-inspired uncertainty adaptation to take into account the distribution shift in uncertainty modeling (subsection 4.2).
@@ 2 Related work
Implicit shape modeling from sparse input
@@
Much attention has recently been directed towards the problem of novel view synthesis from sparse views enhancing this task using NeRFs [8, 23, 24, 25, 26, 27, 28, 29, 30], SDFs [31, 32], and occupancy networks [33]. Implicit functions also found their applications in medical imaging, particularly in 3D shape reconstruction from sparse MRI slices [6, 34, 35, 36, 37, 38]. In this paper, we focus on occupancy networks due to their relevance in medical imaging applications but note that Dropsembles can be used with other neural implicit functions.
Uncertainty modeling
@@
Estimating uncertainty in deep models has attracted significant interest in recent years [17, 18, 39]. A remaining challenge here is uncertainty estimation under distribution shifts, which in most real-world scenarios is essential for guaranteeing the reliability and resilience of predictions when confronted with out-of-distribution (OOD) examples [40, 41, 42]. In 3D reconstruction, this is important since corruption and sparsity pattern in test samples often deviates from those in training data. While numerous methods have been devised to detect OOD cases [43, 44, 45, 46, 47] or enhance accuracy in unobserved domains [48, 49], the adaptation of uncertainty estimates has been relatively unexplored. In a comprehensive benchmark study, [42] find that from all metrics considered, calibration transfers worst, meaning that models that are well calibrated on the training data are not necessarily well calibrated on OOD data. In an era dominated by LLMs and foundation models, addressing this issue becomes paramount, particularly given the tendency of such models for overconfidence as a result of fine-tuning [50, 51]. To overcome such issues, [52] introduces a method for estimating uncertainty in fine-tuned LLMs using Low-Rank Adaptation. In computer vision, [53] highlights that fine-tuning comes at the cost of robustness, and address this issues by ensembling the weights of the zero-shot and fine-tuned models. [54] propose uncertainty estimation in one-shot object detection, particularly focusing on fine-tuned models.
@@
Closer to Dropsembels are [55, 12, 56] that address uncertainty of neural implicit representations. [55] models uncertainty in the color and density output of a scene-level neural representation with conditional normalizing flows. [12] uses this method in robotics applications. [56] estimates uncertainty for neural object representation from monocular images by propagating it from image space first to latent space, and consequently to 3D object shape. Dropsembels can complement all of the above approaches by modeling the uncertainty in the weights of implicit functions throughout fine-tuning, rendering them more robust to distribution shifts. By doing so, it improves both the performance and reliability of methods building on neural implicit representations.
@@ 3.1 Implicit Shape Representations
Implicit functions are widely used in computer vision to represent complex shapes. Assuming a continuous encoding of the object domain $\mathcal{X}$, an SDF for a surface $S$ encoding a shape is defined as $f:\mathcal{X}\to\mathbb{R}$ where: $f(x)<0\Leftrightarrow x\text{ "inside" }S\quad f(x)=0\Leftrightarrow x\in S$. Another popular way to represent a shape object is an occupancy function $g:\mathcal{X}\to\{0,1\}$ defined as $g(x)=\mathbb{I}_{\text{Shape Object}}(x)$ where $\mathbb{I}$ corresponds to the indicator function. The primary advantage of both approaches lies in their ability to accurately model complex shapes from sparse observations using deep neural networks. This capability underpins methods such as "DeepSDF" [57] and "Occupancy Network" [58], which approximate functions by non-linearly regressing observations to the surface encoding the shape. To facilitate learning from training sets, these techniques are often employed in conjunction with latent representations to capture diverse shape variations and at inference time yield continuous shape representations. The latent representations can be trained directly from data using autoencoders [58] or autodecoders [57]. Recent research has also explored using large pre-trained models like CLIP to obtain latent representations [59]. While the specific choice of the latent encoder is beyond the scope of this paper, we will refer to it as an "oracle."
Both methods can be described with the same formulation. We consider a dataset $\mathcal{D}:=\bigg{\{}\{x^{i}_{j},y^{i}_{j}\}_{j\in[K]},Z^{i}\bigg{\}}_{i\in[N]}$, where $N$ is the number of images, $K$ is the number of available pixels per image , $Z^{i}$ is a latent representation of image $i$ given by the "oracle", $x^{i}_{j}\in\mathcal{X}\subseteq\mathbb{R}^{3}$ is the position corresponding to the observed value $y^{i}_{j}$, which are obtained through an SDF or voxel occupancy following the definitions above. In both settings, the regression function is a neural network $f_{\theta}(x,Z),\theta\in\Theta_{NN}$. At training, networks are trained using the following optimization objective.
$\hat{R}_{\mathcal{D}}(\theta)=\frac{1}{NK}\sum_{j\in[K]\ \&\ i\in[N]}l(f_{\theta}(x^{i}_{j},Z^{i}),y^{i}_{j})$ (1)
@@ 3.2 Uncertainty modeling
Deep Ensembles[18] apply ensemble methods to neural networks [60, 61, 62, 63]. Multiple independent networks are trained to determine a set of optimal weights $\{\hat{\theta}_{i}\}_{i\in[M]}$ with the subscript $i$ denoting different networks. Ensembles mitigate the risk of selecting a single set of weights, which may not yield good results particularly when training data consist of fewer samples relative to the size of the parameter space. Instead, multiple solutions with comparable accuracy are determined, allowing the ensemble to average outputs and minimize the selection risk. Training of deep ensembles is associated with high computational demands due to training multiple networks.
@@
Dropout [64] is a regularization technique that aims to reduce overfitting by randomly omitting subsets of features during each training step. By "dropping out" (i.e., setting to zero) a subset of activations within a network layer, it diminishes the network’s reliance on specific neurons, encouraging the development of more robust features. Gal and Ghahramani [17] demonstrated that dropout can also be interpreted from a Bayesian perspective and applied towards modeling uncertainty. Dropout at test time (referred to as Monte Carlo Dropout) is shown to perform approximate Bayesian inference, essentially through using randomness in dropout configurations for uncertainty modeling.
@@ 3.3 Elastic Weight Consolidation
Elastic Weight Consolidation (EWC) is a regularization technique introduced in continual learning to address catastrophic forgetting [22]. The underlying principle is to protect parameters crucial for previous tasks while learning new ones. Assume two distinct tasks, $A$ and $B$, with their respective datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, where $\mathcal{D}_{A}\cap\mathcal{D}_{B}=\emptyset$. The tasks are learned sequentially without access to previous tasks’ datasets. First, task $A$ is learned by training a neural network on $\mathcal{D}_{A}$, resulting in a set of optimal weights $\hat{\theta}_{A}$. When learning task $B$ using dataset $\mathcal{D}_{B}$, EWC regularizes the weights so they remain within a region in the parameter space that led to good accuracy for task $A$.
@@ 4 Methods
Assume access to datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, defined in Section (3.1), with the number of points per image $K_{A}$ and $K_{B}$, and the number of images $N_{A}$ and $N_{B}$. The dataset $\mathcal{D}_{A}$ is assumed to be high quality and "dense," while the dataset $\mathcal{D}_{B}$ is "sparse" in terms of the number of points observed per image and potentially contains corruptions in individual images, i.e., $K_{A}\gg K_{B}$. Additionally, it is assumed that the number of images in $\mathcal{D}_{A}$ is greater than in $\mathcal{D}_{B}$, i.e., $N_{A}>N_{B}$ (in our experiments, we consider the case of a single image $N_{B}=1$). The two datasets can only be accessed successively and not simultaneously; that is, $\mathcal{D}_{A}$ is accessed first, followed by $\mathcal{D}_{B}$, without further access to $\mathcal{D}_{A}$. Note that the latent representation vectors ${Z^{i}}$ are assumed to be obtained by the same learning oracle for both datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ described in Section 3.1. Without loss of generality, in this paper, we consider the latent map $\mathcal{L}:\mathcal{X}\to\mathcal{Z}$ to be a "frozen" encoder pretrained on dataset $\mathcal{D}_{A}$.

### v2

@@ 1 Introduction
Recent advancements in neural implicit functions have facilitated their use for 3D object representations and applications in novel views synthesis in computer vision. Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) have gained recognition for their ability to accurately synthesize novel views of complex scenes and became an important tool in applications of photorealistic rendering (Martin-Brualla et al., 2021), such as virtual reality and augmented reality. Signed Distance Functions (SDFs) (Park et al., 2019a; Gropp et al., 2020) is shown to be particularly useful in scenarios where precise boundary details are crucial, like industrial design and robotics. Occupancy Networks (Mescheder et al., 2019b), which model shapes as a probabilistic grid of space occupancy, excel in handling topological variations, making them ideal for medical imaging (Amiranashvili et al., 2022) and animation. Advancing this concept, Convolutional Occupancy Networks (Peng et al., 2020) integrate convolutional networks to enhance spatial learning, proving effective in detailed architectural modeling and complex reconstructions.
@@
In practice, achieving optimal performance with these methods often requires densely sampled input data and a high degree of similarity between the training data and the target object. However, such ideal conditions are rarely met in practical applications like augmented reality, virtual reality, autonomous driving, and medical contexts, where inputs are typically sparse and less precise (Truong et al., 2023; Müller et al., 2022a). For example, in medical applications, generating precise 3D representations from sparse inputs is particularly crucial for morphological analysis (Tóthová et al., 2020). Patient-specific anatomical modeling significantly enhances the assessment of a patient’s condition and aids in devising customized treatment plans (Turella et al., 2021).
@@
In addition to sparsity, real-world data usually suffers from noise and corruption that leads to distribution shifts with respect to the training data, and thus to significant performance degradation. Occlusion, noise, truncation, and lack of depth measurements (Liao et al., 2023) greatly affect reconstruction from monocular observations. Noise in estimations of camera poses inevitably degrades reconstruction from sparse views (Truong et al., 2023; Zhang et al., 2022; Chen et al., 2021b). Under input sparsity and distribution shifts, in safety-critical applications, such as medical (Zou et al., 2023) or autonomous driving (Kiran et al., 2021), it is crucial that inferred information is transparently disclosed to the end user, as it may significantly influence the decision-making process. One approach to this end is to quantify uncertainty in the reconstruction. For example, Shen et al. (2021) derived a variational inference objective to model predictive distribution in NeRFs. However, similar results are yet to be established for more generic applications of implicit functions.
To tackle the challenges of sparse and corrupted inputs, we propose to leverage dense, noise-free synthetic data for pre-training. For instance, a pretrained encoder can be shared across both synthetic and real datasets, while a lightweight implicit decoder is initially trained on synthetic data and subsequently fine-tuned on the sparse, noisy real data. This process introduces two primary sources of uncertainty: from the encoder’s latent representation and the fine-tuning of the decoder. In this paper, we focus on the latter and demonstrate that epistemic uncertainty is an important, yet understudied source of error in implicit functions, which significantly impacts the predictions quality. To the best of our knowledge, this type of uncertainty quantification in 3D reconstruction for fine-tuned neural implicit functions has not been addressed in the literature.
@@
Monte Carlo (MC) dropout (Gal and Ghahramani, 2016) and deep ensembles (Lakshminarayanan et al., 2017) are two commonly used baselines for estimating uncertainty in computer vision applications (Kendall and Gal, 2017), which could be readily applied to our task. MC dropout is simple to integrate and computationally efficient during training. However, it often underestimates uncertainty and requires multiple forward passes during inference, increasing computational cost (Postels et al., ; Postels et al., 2021). Deep ensembles involve training multiple neural networks independently and averaging their predictions, capturing a wider range of potential outputs. This method provides improved predictive performance and better-calibrated uncertainty estimates but is computationally costly and memory-intensive. This becomes especially demanding in our fine-tuning context, where each model in an ensemble needs to be trained on both datasets. Here we introduce Dropsembles [^1] https://github.com/klanita/Dropsembles, a method that creates ensembles based on the dropout technique. Dropsembles aim to moderate the computational demands associated with ensembles while attempting to maintain prediction accuracy and uncertainty calibration of deep ensembles. Combined with Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2016), Dropsembles is able to mitigate distribution shifts between source and target datasets.
@@
  • To the best of our knowledge, we are the first to model epistemic uncertainty in the implicit decoder during fine-tuning from synthetic data.
  • To this end, we introduce Dropsembles (overview in Figure 1) to achieve the performance of vanilla ensembles with significantly reduced computation cost (subsection 4.1).
  • We introduce EWC-inspired regularization for task-agnostic uncertainty adaptation to take into account the distribution shift in uncertainty modeling (subsection 4.2).
@@ 2 Related work
Implicit shape modeling from sparse input Much attention has recently been directed towards the problem of novel view synthesis from sparse views enhancing this task using NeRFs (Truong et al., 2023; Chen et al., 2021a; Deng et al., 2022; Jain et al., 2021; Kim et al., 2022; Niemeyer et al., 2022; Roessle et al., 2022; Lin et al., 2021; Oechsle et al., 2021), SDFs (Yu et al., 2022; Yariv et al., 2020), and occupancy networks (Niemeyer et al., 2020). Implicit functions also found their applications in medical imaging, particularly in 3D shape reconstruction from sparse MRI slices (Amiranashvili et al., 2022; 2024; Gatti et al., 2024; McGinnis et al., 2023; Alblas et al., 2023; Jacob et al., 2023). In this paper, we focus on occupancy networks due to their relevance in medical imaging applications but note that Dropsembles can be used with other neural implicit functions.
@@
Uncertainty modeling Estimating uncertainty in deep models has attracted significant interest in recent years (Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Gawlikowski et al., 2023). A remaining challenge here is uncertainty estimation under distribution shifts, which in most real-world scenarios is essential for guaranteeing the reliability and resilience of predictions when confronted with out-of-distribution (OOD) examples (Tran et al., 2022; Li et al., 2022; Wenzel et al., 2022). In 3D reconstruction, this is important since corruption and sparsity pattern in test samples often deviates from those in training data. While numerous methods have been devised to detect OOD cases (Malinin and Gales, 2018; Tagasovska and Lopez-Paz, 2019; Ren et al., 2019; Charpentier et al., 2020; Osband et al., 2023) or enhance accuracy in unobserved domains (Liu et al., 2021; Hendrycks et al., 2021), the adaptation of uncertainty estimates has been relatively unexplored. In a comprehensive benchmark study, Wenzel et al. (2022) find that from all metrics considered, calibration transfers worst, meaning that models that are well calibrated on the training data are not necessarily well calibrated on OOD data. In an era dominated by LLMs and foundation models, addressing this issue becomes paramount, particularly given the tendency of such models for overconfidence as a result of fine-tuning (Guo et al., 2017a; Dodge et al., 2020). To overcome such issues, (Balabanov and Linander, 2024) introduces a method for estimating uncertainty in fine-tuned LLMs using Low-Rank Adaptation. In computer vision, (Wortsman et al., 2022) highlights that fine-tuning comes at the cost of robustness, and address this issues by ensembling the weights of the zero-shot and fine-tuned models. (Lu and Koniusz, 2022) propose uncertainty estimation in one-shot object detection, particularly focusing on fine-tuned models.
@@
Several works focused on modeling uncertainty in the encoder of neural implicit representations. (Shen et al., 2022) models uncertainty in the color and density output of a scene-level neural representation with conditional normalizing flows. Liao et al. (2023) uses this method in robotics applications. Liao and Waslander (2024) estimates uncertainty for neural object representation from monocular images by propagating it from image space first to latent space, and consequently to 3D object shape. Yang et al. (2023) focused on modeling aleatoric uncertainty in occupancy networks. Dropsembles can complement all of the above approaches by modeling the epistemic uncertainty in the weights of implicit functions throughout fine-tuning, rendering them more robust to distribution shifts. By doing so, it improves both the performance and reliability of methods building on neural implicit representations.
@@ 3.1 Implicit Shape Representations
Implicit functions are widely used in computer vision to represent complex shapes. Assuming a continuous encoding of the object domain $\mathcal{X}$, an SDF for a surface $S$ encoding a shape is defined as $f:\mathcal{X}\to\mathbb{R}$ where: $f(x)<0\Leftrightarrow x\text{ ``inside'' }S\quad f(x)=0\Leftrightarrow x\in S$. Another popular way to represent a shape object is an occupancy function $g:\mathcal{X}\to\{0,1\}$ defined as $g(x)=\mathbb{I}_{\text{Shape Object}}(x)$ where $\mathbb{I}$ corresponds to the indicator function. The primary advantage of both approaches lies in their ability to accurately model complex shapes from sparse observations using deep neural networks. This capability underpins methods such as “DeepSDF” (Park et al., 2019b) and “Occupancy Network” (Mescheder et al., 2019a), which approximate functions by non-linearly regressing observations to the surface encoding the shape. To facilitate learning from training sets, these techniques are often employed in conjunction with latent representations to capture diverse shape variations and at inference time yield continuous shape representations. The latent representations can be trained directly from data using autoencoders (Mescheder et al., 2019a) or autodecoders (Park et al., 2019b). Recent research has also explored using large pretrained models like CLIP to obtain latent representations (Cheng et al., 2023). While the specific choice of the latent encoder is beyond the scope of this paper, we will refer to it as an “oracle”.
Both methods can be described with the same formulation. We consider a dataset $\mathcal{D}:=\bigg{\{}\{x^{i}_{j},y^{i}_{j}\}_{j\in[K]},Z^{i}\bigg{\}}_{i\in[N]}$, where $N$ is the number of images, $K$ is the number of available pixels per image , $Z^{i}$ is a latent representation of image $i$ given by the “oracle”, $x^{i}_{j}\in\mathcal{X}\subseteq\mathbb{R}^{3}$ is the position corresponding to the observed value $y^{i}_{j}$, which are obtained through an SDF or voxel occupancy following the definitions above. In both settings, the regression function is a neural network $f_{\theta}(x,Z),\theta\in\Theta_{NN}$. At training, networks are trained using the following optimization objective.
$\hat{R}_{\mathcal{D}}(\theta)=\frac{1}{N\cdot K}\sum_{j\in[K]\ \wedge\ i\in[N]}l(f_{\theta}(x^{i}_{j},Z^{i}),y^{i}_{j})$ (1)
@@ 3.2 Uncertainty modeling
Deep Ensembles (Lakshminarayanan et al., 2017) apply ensemble methods to neural networks (Lee et al., 2020; Breiman, 2001; Schapire, 1990; Saberian and Vasconcelos, 2010). Multiple independent networks are trained to determine a set of optimal weights $\{\hat{\theta}_{i}\}_{i\in[M]}$ with the subscript $i$ denoting different networks. Ensembles mitigate the risk of selecting a single set of weights, which may not yield good results particularly when training data consist of fewer samples relative to the size of the parameter space. Instead, multiple solutions with comparable accuracy are determined, allowing the ensemble to average outputs and minimize the selection risk. Training of deep ensembles is associated with high computational demands due to training multiple networks.
@@
Dropout (Srivastava et al., 2014) is a regularization technique that aims to reduce overfitting by randomly omitting subsets of features during each training step. By “dropping out” (i.e., setting to zero) a subset of activations within a network layer, it diminishes the network’s reliance on specific neurons, encouraging the development of more robust features. Gal and Ghahramani (Gal and Ghahramani, 2016) demonstrated that dropout can also be interpreted from a Bayesian perspective and applied toward modeling uncertainty. Dropout at test time (referred to as Monte Carlo Dropout) is shown to perform approximate Bayesian inference, essentially through using randomness in dropout configurations for uncertainty modeling.
@@ 3.3 Elastic Weight Consolidation
Elastic Weight Consolidation (EWC) is a regularization technique introduced in continual learning to address catastrophic forgetting (Kirkpatrick et al., 2016). The underlying principle is to protect parameters crucial for previous tasks while learning new ones. Assume two distinct tasks, $A$ and $B$, with their respective datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, where $\mathcal{D}_{A}\cap\mathcal{D}_{B}=\emptyset$. The tasks are learned sequentially without access to previous tasks’ datasets. First, task $A$ is learned by training a neural network on $\mathcal{D}_{A}$, resulting in a set of optimal weights $\hat{\theta}_{A}$. When learning task $B$ using dataset $\mathcal{D}_{B}$, EWC regularizes the weights so they remain within a region in the parameter space that led to good accuracy for task $A$.
@@ 4 Methods
Assume access to datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$, defined in Section 3.1, with the number of points per image $K_{A}$ and $K_{B}$, and the number of images $N_{A}$ and $N_{B}$. The dataset $\mathcal{D}_{A}$ is assumed to be high quality and “dense”, while the dataset $\mathcal{D}_{B}$ is “sparse” in terms of the number of points observed per image and potentially contains corruptions in individual images, i.e., $K_{A}\gg K_{B}$. Additionally, it is assumed that the number of images in $\mathcal{D}_{A}$ is greater than in $\mathcal{D}_{B}$, i.e., $N_{A}>N_{B}$ (in our experiments, we consider the case of a single image $N_{B}=1$). The two datasets can only be accessed successively and not simultaneously; that is, $\mathcal{D}_{A}$ is accessed first, followed by $\mathcal{D}_{B}$, without further access to $\mathcal{D}_{A}$. Note that the latent representation vectors ${Z^{i}}$ are assumed to be obtained by the same learning oracle for both datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ described in Section 3.1. Without loss of generality, in this paper, we consider the latent map $\mathcal{L}:\mathcal{X}\to\mathcal{Z}$ to be a “frozen” encoder pretrained on dataset $\mathcal{D}_{A}$.